If f(x)=cos(x) - sin(x) then what is the value of f(1)?

Grade:11

2 Answers

Dev kumar basant
6 years ago
anwer will be negative and in fraction you can see easily in the graph of sinx and cosx they intersect at 45
and after that sinx is above then cosx.
Arun
6 years ago
here in f(1), 1 is angle in radian.
so f(1)= cos1−sin1
this will be a negative answer as in [0,\pi/4) cos will be greater than sin.
and 1 will be less than \pi/4.